IF lea is 6 years older than Sue and john is 5 yrs older than lea and the toatl ages are 41 then how old is Sue.
Got question from teas test preview website

Answer by SHUgrad05(58)

Let Sue's age=x
Lea is 6 years older than Sue= x+6
John is 5 years older than Lea= (x+6)+5=x+11
Their total ages=41
x+x+6+x+11=41
3x+17=41
3x=24
x=8
Sue is 8
Lea is 14
John is 19